A thick cylinder of 100 mm external diameter and 50 mm internal diameter is wound with

steel wire of 1 mm diameter, initially stressed to 20 MN/m’ until the outside diameter is

120 mm. Determine the maximum hoop stress set up in the cylinder if an internal pressure of

30 MN/m’ is now applied.


1
Expert's answer
2021-08-27T08:30:10-0400

Given,

External diameter "(R)=100 mm=100\times 10^{-3}m"

Internal diameter "(r)= 50 mm=50\times 10^{-3}m"

"r_1=0.025"

Diameter of steel wire "(d)=1mm=10^{-3}m"

Applied stress"= 20MN/m =20\times 10^6N/m"

Outside diameter "(d)=120mm =120\times 10^{-3}m"

"(r_2)=0.06m"

Internal pressure "(\sigma_r)=30MN/m = 30\times 10^6N/m"


"\Rightarrow -30=A-\frac{B}{0.025}=A-1600B...(i)"


"\Rightarrow 0=A-\frac{B}{0.06^2}=A-278B ...(ii)"

From equation (i) and (ii),

"\Rightarrow 30=1322B"


"\Rightarrow B=\frac{30}{1322}=0.0227"


"\Rightarrow A=278B"

"\Rightarrow A=278\times 0.0227 =6.32"


For calculating the stress, the wire and cylinder should be treated like a single thick cylinder with internal diameter 50mm and 120mm external diameter.

Hoop stress at 25mm is "(\sigma_{r_1})=A+1600B=42.7MN/m^2"


Hoop stress at 60mm is "(\sigma_{r_2})=15.4MN/m^2"


"\Rightarrow \sigma_r=-\frac{r^2-R^2}{2r^2}T\log(\frac{R_3^2-R_1^2}{r^2-R_1^2})=p"


"\Rightarrow \sigma_r=\frac{0.05^2-0.025^2}{2\times 0.05^2}T\log_e(\frac{0.06^2-0.025^2}{0.05^2-0.025^2})"


"\Rightarrow \sigma_r = \frac{18.75\times 20}{50}\log_e(\frac{29.75}{18.75})"


"\Rightarrow \sigma_r=-3.45MN/m^2"

So, for the wire winding, the stress in the tube

"\sigma_r=-3.45" at "r=0.05"

and "\sigma_r =0" to "r=0.025"

"\sigma _r=A-400B"

"\Rightarrow -3.45=A-400B ....(iii)"

and "0=A-1600B...(iv)"

Now, solving equation (iii) and equation (iv)

"-3.45=1200B"

"\Rightarrow B=2.88\times 10^{-3}" and "A=-4.6"

Hoop stress at 25mm is "(\sigma_{r_1})=A+1600B=-9.2MN/m^2"

Hoop stress at 50mm is "=A+400B =-5.75 MN\/m^2"

Hence, the internal pressure at

r=25mm, "\sigma_H=42.7-9.2=33.5MN/m^2"

"r=50mm," "\sigma_H=15.4-5.75 =9.65MN/m^2"

So, the maximum hoop stress "33.5MN\/m^2"